2. Let the equation f(x) = 1.2 ln x + 2.3 represent the wind speed in meters per second at a height x meters above the ground.
a. Find the wind speed at 4 meters above ground.
b. Estimate the height needed to achieve a wind speed of 6 meters per second.

Answer by alicealc(293)

(a) wind speed at 4 meters above the ground:
f(4) = 1.2*ln(4) + 2.3 = 1.2*1.3863 + 2.3 = 1.6636 + 2.3 = 2.9636 meters per second

(b) height needed for wind speed of 6 meters per second:
6 = 1.2*ln(x) + 2.3
6 - 2.3 = 1.2*ln(x)
3.7 = 1.2*ln(x)
3.7/1.2 = ln(x)
3.0833 = ln(x)
x = e^3.0833 = 21.8303 meters